OpenStudy (anonymous):

-3y + 5 (6y+9) = -1 -5 (4-y) Solve for y

OpenStudy (anonymous):

this is the problem

jimthompson5910 (jim_thompson5910):

-3y+5(6y+9)=-1-5(4-y) -3y+30y+45=-1-20+5y 27y+45=-1-20+5y 27y+45=5y-21 27y=5y-21-45 27y-5y=-21-45 22y=-21-45 22y=-66 y=-66/(22) y=-3

OpenStudy (anonymous):

i've tried -3.. its incorrect

OpenStudy (anonymous):

WTF! NVM!!!

OpenStudy (anonymous):

I just reput it down, and it said its correct! UGH! Thanks!!!!

jimthompson5910 (jim_thompson5910):

lol weird, but ok

jimthompson5910 (jim_thompson5910):

i saw -33 for your initial answer, that's probably what you typed at first